I am an old student back in college taking Beginning algebra. We are doing systems of equations and inequalities. Two questions involving solving eq. by addition. 4(x+1)=17-3(y-1),2(x+2)+3(y-1)=9
Second question:3x-16=5y,33-5y=4x both are systems.
Thank you,Liz

Answer by mukhopadhyay(490)

4(x+1)=17-3(y-1)
=>4x+4=17-3y+3
=>4x+3y=17+3-4
=>4x+3y=16
2(x+2)+3(y-1)=9
=>2x+4+3y-3=9
=>2x+3y=9-4+3
=>2x+3y=8
I would use the first equation to get 3y interms of x, so that I can substitute it to the second equation
4x+3y=16 =>3y=16-4x
Second Equation, 2x+3y=8
=>2x+(16-4x)=8 (I substituted)
=>2x+16-4x=8
=>2x-4x=8-16
=>-2x=-8
=>x=4
Using the value of x=4 on 3y=16-4x,
3y=16-16
=>y=0
The answer is, x=4 and y=0